OpenStudy (anonymous):

-x + 3y = 18 -6x – 3y = 3 help me plx

ganeshie8 (ganeshie8):

add both the equations vertically so that \(y\) stuff cancels out

ganeshie8 (ganeshie8):

-x + 3y = 18 -6x – 3y = 3 -------------------

ganeshie8 (ganeshie8):

add them, wat do u get ?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

Add downward. The -x +-6x, the 3y+ the -3y and the 18 +3 this will give you a new equation.

OpenStudy (anonymous):

?x+?y=?

OpenStudy (anonymous):

I'll give you a hint.\[-7x+?y=?\]

OpenStudy (anonymous):

i need 2 answers

ganeshie8 (ganeshie8):

-x + 3y = 18 -6x – 3y = 3 ------------------- -7x + 0 = 21 -7x = 21 x = -3

ganeshie8 (ganeshie8):

you got one answer, x = -3

ganeshie8 (ganeshie8):

simply substitute x = -3 in first equation, and solve for y

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

-6x-3y=3

OpenStudy (anonymous):

-6x-3y=3

ganeshie8 (ganeshie8):

we got x = -3 simply put this value in first equation, and solve for y -x + 3y = 18 -(-3) + 3y = 18 3 + 3y = 18 3y = 15 y = 5

ganeshie8 (ganeshie8):

so the solution to given system is : x = -3, y = 5

ganeshie8 (ganeshie8):

see if that makes some sense :)

OpenStudy (anonymous):

thank you